the product of two consecutive positive intergers is 19 more than their sum.find the integers...


thank you

Answer by mananth(16946)

let the integers be x,x+1
x(x+1)=x+x+1+19
x^2+x=2x+20
x^2-x-20=0
(x-5)(x+4)=0
x= 5 OR -4
the numbers are
5,6 and -4,-3
...
m.ananth@hotmail.ca
...